The components of vectors B? and C? are given as follows:

The angle (less than 180 degrees) between vectors  B? and C?, in degrees, is closest to:

The components of vectors B? and C?
Answer #1

Based on the given components,

As B has both negative components, It lies in 3rd quadrant

And C has a positive Y and negative x component. So, it lies in 2nd quadrant

Angle made by B with x - axis = arctan(By / Bx) = arctan(-7.4 / -7.3) = 45.39 degrees below the x - axis in third quadrant

Angle made by C with x - axis = arctan(Cy / Cx) = arctan(3 / -8.1) = 159.68 with positive x - axis

That is C makes (180 - 159.68) = 20.323 with negative x - axis

So, the angle between B and C will be = 20.323 + 45.39 = 65.71 or approximately 66 degrees